# What is the nature of Little Red Dots and what is not, MIRI SMILES edition

### v1

@@ 5.3 Physical properties of the LRDs detected in the bluest MIRI bands
All other LRDs in our sample detected up to F1000W, apart from the Golden Five, are shown jointly in the SED plots provided in Figure 11. In the same way, the SED fits for all F770W detections not included in any previous SED plot, as well as all sources not detected by MIRI are shown and discussed in Appendix C. Physical properties derived from each code for all individual galaxy as well as statistical properties obtained for the whole sample and subsamples are given in Tables 2 and 3.
@@
Even though the dust emission spectral region is not fully probed for the 10 m sample, and to a much larger extent the 7.7 m and non-MIRI samples, the upper limits imposed by the MIRI data at longer wavelengths, more specifically, at 12.8 and 15 m, indicate a very similar behavior of the spectral range around 2-3 m compared to what we showed for the Golden Five galaxies. Indeed, the SED flattens, indicating that possible dust emission powered by stellar or AGN heating is not dominant and could only start adding significant flux redward of 3 m. The properties we infer from the UV-to-NIR SEDs are also similar for the F1000W, F770W, and no-MIRI subsamples compared to those obtained for the Golden Five galaxies. There are, however, some observational trends, which translate to differences in physical properties.
@@
First, the Golden Five galaxies are brighter than the rest of the sources in the full sample (cf. Figure 1). The median and quartiles for F444W are mag, compared to mag for the F1000W sample, mag for the F770W sample, and for the sources with no MIRI detection. Concerning colors, the F1000W sample is redder than the Golden Five, and the overall sample. As shown in Figures 1 and 4, the F1000W sample is among the reddest in both the NIRCam and MIRI colors: F277W-F444W=2.2 mag, F444W-F770W=1 mag and F770W-F1000W=0.9 mag, versus the overall sample, 1.4 mag, 0.7 mag and 0.2 mag respectively. Comparatively, the largest difference is in F770W-F1000W, where the F1000W sample has similar colors to 2 galaxies among the Golden Five, JADES-79803 and JADES-211388, whose MIR spectral range is fitted with dust tori. We note that for JADES-211388, at , F1000W lies on top of the He IPa- line, which partially explains the red color.
@@
Looking at rest-frame colors and stacked SEDs in Figures 10 and 11, we find that the F1000W sample has also redder 0.4-to-1.0 m colors than the Golden Five sources, suggesting that they are dustier (see next paragraphs). Interestingly, using a longer baseline color 0.25-to-1.0 m (similar to NUV-J), which probes into the relatively flat UV SED of the LRDs, we find even larger differences between the Golden Five galaxies and the F1000W samples, 0.25-to-1.0 m2.1 mag vs. 3.5 mag. These colors and the multi-component SED modeling discussed in the previous section lead to conclude that the colors relative to rest 1 m are partially driven by differences in the relative luminosity of the component dominating the rest-UV (young, unobscured stellar population) and the component dominating the rest-optical (older stellar population or obscured accretion disk). As the flat UV component scales up in brightness, it leads to bluer 0.25-to-1.0 m and 0.4-to-1.0 m colors, but perhaps not because of a change in the intrinsic properties of the component dominating the optical range. It could be that a brighter UV reveals a larger fraction of the starburst emission percolating through the compact dust (possibly linked to a higher burst strength or younger age, apart from dust-star relative geometry), or perhaps it shows a more massive stellar host for the obscured AGN. This interpretation also helps to explain the larger scatter in the UV region of the stacked SEDs relative to the optical region. I.e., while all LRDs have distinctive blue-UV and red-optical SEDs, there is a larger diversity in the UV emission for a similar optical-to-NIR slope that might reflect variations in the relative luminosity of two different components.
@@
The overall colors of the LRDs range from 0.25-to-1.0 m2 to 4.5 mag. Taking 0.25-to-1 m3 mag as an intermediate value, we find that all of the Golden Five galaxies exhibit bluer colors, versus only 30% (2/7) of the galaxies in the F1000W and F770W samples. This emphasizes again that the Golden Five galaxies are intrinsically bluer than the other samples in all colors. Interestingly, we also find a trend toward stronger emission lines (larger EWs) with bluer 0.25-to-1.0 m colors. For example, among the bluest Golden Five sources (JADES79803 and JADES204851) and in the handful of galaxies with bluer colors in the F1000W (JADES210600, JADES214552 and JADES217926) and F770W samples (JADES187025 and JADES197348).
@@
Concerning stellar masses, the Golden Five galaxies are 0.2-0.4 dex larger than the other subsamples, with values around for the former. Interestingly, the stellar population attenuation is lower for the Golden Five galaxies, mag, compared to F1000W sources ( mag), F770W galaxies ( mag), and non-MIRI sources ( mag).
@@ 5.4 Statistical stellar and AGN properties of LRDs
The typical stellar mass of LRDs is (median and quartiles) according to synthesizer-AGN. Considering the full redshift range of our sample of LRDs, , and the number density of galaxies in the stellar mass range detected by CEERS based on the v0.51 catalogs (Finkelstein et al., 2023), we calculate that the number of LRDs account for % of the full population of galaxies. This translates to a comoving density of LRDs of Mpc, which is quite constant across the redshift range, with differences dex between and . The estimates for both the LRDs and the other galaxies are subject to a number of potential systematic errors, but the estimate indicates that the LRDs represent a significant, but not a dominant, population over this redshift range. Given the Universe ages probed by our sample, 200 Myr for and 400 Myr for , the % frequency within the global population, could be interpreted in terms of a duty cycle around 20–40 Myr, which point to a starburst phase.
Larger masses are obtained by prospector-SF, , mainly because the mass-weighted ages are older, between 10 and 100 Myr, with median and quartiles being Myr. synthesizer-AGN fits the SEDs with significantly younger stellar populations, typically Myr. The 2 stellar populations considered by synthesizer-AGN are typically younger than 20 Myr. In fact, the average SFH of LRDs obtained by synthesizer-AGN and shown in Figure 13 indicates that LRDs are experiencing a very intense episode of star formation extending nearly 10 Myr and with a very compact size. The burst would be in part heavily dust-enshrouded, with some younger stars having cleared the interstellar medium and being directly observable through much smaller dust optical paths. The young ages are expected for starbursts with strong emission lines (as present in some of our galaxies), with large amounts of dust (confirmed for a significant fraction of the whole sample), and with gas also feeding a SMBH. synthesizer-SF provides a similar average SFH, extended almost at a constant level up to approximately 10 Myr, and decaying afterwards. However, the first age bin considered by synthesizer-SF encompasses the 2 bursts obtained by synthesizer-AGN, the former adding more mass in ages around 10 Myr (with a large scatter, shown by the shaded region in Figure 13).
@@
The bolometric luminosities (including dust-absorbed energy) vary by a factor of 10 between the prospector-SF and prospector-AGN runs, while synthesizer-AGN lie in between, with a typical value of . Combining with the stellar masses, we conclude that LRDs present mass-to-light ratios of 1/400, typical of OB stellar associations (a B2 star would have that value, approximately), as expected based on the young ages. We would like to remark here that given these very young ages and the starburst nature of LRDs, the a priori assumption of a universal IMF is quite relevant. The amount of OB stars formed, the quick and efficient formation of metals and dust, and the inferred stellar masses (or even the growth of a SMBH) are all affected by the IMF.
@@ 5.5 Spectroscopic properties of LRDs found in the literature
Apart from JADES204851, whose broad-band emission of the H emission at (2000 km s), published in Matthee et al. (2023b) was mentioned in Section 5, there are other sources in our sample with relevant spectroscopic information. JADES197348 was included in the JADES NIRSpec initial data release (Bunker et al., 2023) and identified with a broad-line AGN (Maiolino et al., 2023b). Its spectrum shows a km s component that accounts for two thirds of the H total flux, while [OIII] shows no such component. Our fits to this source are dominated by a QSO-like spectrum in the optical and NIR in the case of synthesizer-AGN and prospector-AGN. JADES154428 is found to present a broad-line component with FWHM 1800 km s (Sun et al. 2024, in prep.), our fits include a non-negligible contribution from a QSO-like spectrum, dominating the SED (prospector-AGN and prospector-AGN or accounting for 50% of the emission at specific wavelengths (synthesizer-AGN). No other broad H or H line component has been reported in the rest of 18 galaxies with available spectroscopy (i.e., 17% of the spectroscopic sample are confirmed AGN hosts). We note that the typical depth of FRESCO NIRCam 3.9–-5.0-m grism spectroscopy is erg s cm for broad emission line (FWHM 1000 km s) from a point source and, therefore, faint broad line emission (possibly from an AGN not dominated the continuum) can remain undetected.
@@ 6 Summary and conclusions
We characterize the nature of Little Red Dots (LRDs) in the JADES field by analyzing their spectral energy distributions including the mid-infrared fluxes provided by the SMILES program for all MIRI broad-band filters, which probe the rest-frame near-and mid-infrared where stellar emission and/or obscured AGN emission peak. After removing brown dwarfs, which contaminate our sample of F277W-F444W1 mag, F150W-F200W0.5 mag, F444W28 mag sources at the 15% level, we arrive at a sample of 31 LRDs, the surface density being 0.9 arcmin. This translates to a number density Mpc, accounting for 10–15% of the global population of galaxies with similar redshifts () and stellar mass (). Two thirds are detected in the F560W and F770W filters (all sources brighter than F444W26.5 mag), two fifths in F1000W, one seventh in F1500W, one thirteenth in F1800W, and one source in F2100W, down to 5 limits between 26.1 and 22.6 mag. The MIRI detection fraction is largely dependent on the F444W brightness, but we find an additional trend towards more detections at 10 m or beyond with bluer F277W-F444W colors.
We find that the observed MIRI colors of the LRDs, in combination with the reddest NIRCam bands, are bluer than the typical obscured QSO templates which are dominated by the torus warm/hot dust emission at m. Indeed, the rest-frame NIR spectral range exhibits a much shallower slope which is consistent with the peak of stellar emission at around m.
@@
In general, stellar-dominated models obtain a better agreement to the near-infrared at 1-2 m as well as in the UV at m. The AGN-dominated models where emission from an obscured accretion disk dominates the optical and NIR and the torus takes over at longer wavelengths m provide a better agreement than the typical QSO templates, but still worse than the stellar models. Furthermore, this model also has conceptual problems given that many of the LRDs do not present emission lines, which should be expected for the direct detection of the accretion disk (and its broad- and narrow-line regions). Consequently, we favor the interpretation that the UV-to-optical spectral range of most LRDs is dominated (% luminosity ratio) by stars.
@@
Very young stellar ages (typically around 10 Myr or younger) are supported by high equivalent width lines seen in some LRDs (with spectroscopic and imaging data), and the presence of large amounts of dust, which is a common feature of all models for the full sample, and can be expected for gas-rich, quickly enriched dense starbursts with large amounts of OB stars (affected by the IMF). However, despite the presence of large amounts of dust, LRDs are characterized by a relatively blue emission at wavelengths m, which can only be reproduced if this spectral range is dominated by a QSO, holes in the interstellar medium surrounding star-forming region which allows us to see unobscured very young star formation, or a gray attenuation law typically linked to significant scattering (with implications to the relative geometry of dust and stars and the dust clumpiness). Indeed, the large 3 mag scatter in the UV-to-NIR colors within the typical blue-red SED of the LRDs might indicate different burst strengths of fractions of percolating light through the compact dust envelope. The presence of a QSO is hinted by some medium-band colors consistent with MgII emission. Nevertheless, for most of the sources, the best spectral energy distribution fit is obtained with just stars presenting a differentiated and/or gray dust attenuation law. A significant fraction of the total stellar emission of LRDs comes from OB stellar clumps mostly embedded in dense dusty regions with large optical paths, mag, with the integrated stellar emission obscured in a 90-95%, which would explain the near- and mid-infrared emission jointly with some contribution from an obscured AGN.
@@
| Sample (N) | redshift | code | [mag] | m-w age [Myr] | ||
@@ Appendix B Dust emission modeling
In Figure 16, we show models for a L and L compact (350 pc in size) region, attenuations mag, OB luminosity ratios 40 and 90%, and hydrogen number densities and cm. The templates include a fixed stellar population, which is removed for our modeling of LRDs. Models with and without stars are shown in Figure 16.
@@
The MIR slopes of the Siebenmorgen & Krügel (2007) models are very similar to the slopes of AGN torus emission presented in Nenkova et al. (2008), which are used in our prospector-AGN fits. Nenkova et al. (2008) parametrizes the emission from clumpy dust tori in terms of the number of clouds intercepting the visual, which depend on other parameters such as the torus thickness (outer to inner radii), the density and angular distribution of clumps and the viewing angle. The subset of templates included in FSPS and then Prospector adopts typical assumptions for the majority of these parameters (see e.g., Leja et al. 2018) leaving e only the scaling (overall luminosity) and the optical depth of an individual dust clump at 5500Å( from 5 to 150) as free parameters. For the prospector-AGN fits we remove any contribution from the accretion disk in the dust tori templates and we instead model the disk separately following a combination of empirical QSO template plus a power-law f= (equivalent to f=) with variable slope to 0.5, as discussed in Hernán-Caballero et al. 2016 or more recently in Bosman et al. 2023, attenuated by a Calzetti law. For reference, the accretion disk model with A(V)=1.5 is shown in Figure 16 in magenta. The main difference between the Nenkova et al. (2008) models and the ones in Siebenmorgen & Krügel (2007) dominated by star formation is the absence of PAH emission (especially relevant for our observations, the one at 3 m) and the silicate absorption present in some dust models. Overall, a dust torus (and an AGN template) is almost featureless.

### v2

@@ 5.3 Physical properties of the LRDs detected in the bluest MIRI bands
All other LRDs in our sample detected up to F1000W, apart from the Golden Five, are shown jointly in the SED plots provided in Figure 11. In the same way, the SED fits for all F770W detections not included in any previous SED plot, as well as all sources not detected by MIRI are shown and discussed in Appendix C. Physical properties derived from each code for individual galaxies as well as statistical properties obtained for the whole sample and subsamples are given in Tables 2 and 3.
@@
Even though the dust emission spectral region is not fully probed for the 10 m sample, and to an even lesser extent the 7.7 m and non-MIRI samples, the upper limits imposed by the MIRI data at longer wavelengths, more specifically, at 12.8 and 15 m, indicate a very similar behavior of the spectral range around 2-3 m (rest) compared to what we showed for the Golden Five galaxies. Indeed, the SED flattens, indicating that possible dust emission powered by stellar or AGN heating is not dominant and could only start adding significant flux redward of 3 m. The properties we infer from the UV-to-NIR SEDs are also similar for the F1000W, F770W, and no-MIRI subsamples compared to those obtained for the Golden Five galaxies. There are, however, some observational trends, which translate to differences in physical properties.
@@
First, the Golden Five galaxies are brighter than the rest of the sources in the full sample (cf. Figure 1). The median and quartiles for F444W are mag, compared to mag for the F1000W sample, mag for the F770W sample, and for the sources with no MIRI detection. Concerning colors, the F1000W sample is redder than the Golden Five, and the overall sample. As shown in Figures 1 and 4, the F1000W sample is among the reddest in both the NIRCam and MIRI colors: F277W-F444W=2.2 mag, F444W-F770W=1 mag and F770W-F1000W=0.9 mag, versus for the overall sample they are 1.4 mag, 0.7 mag and 0.2 mag respectively. Comparatively, the largest difference is in F770W-F1000W color difference, where the F1000W sample has similar colors to 2 galaxies among the Golden Five, JADES-79803 and JADES-211388, whose MIR spectral range is fitted with dust tori. We note that for JADES-211388, at , F1000W lies on top of the He IPa- line, which partially explains the red color.
@@
Looking at the rest-frame colors and stacked SEDs in Figures 10 and 11, we find that the F1000W sample has also redder 0.4-to-1.0 m colors than the Golden Five sources, suggesting that they are dustier (see next paragraphs). Interestingly, using a longer baseline color 0.25-to-1.0 m (similar to NUV-J), which probes into the relatively flat UV SED of the LRDs, we find even larger differences between the Golden Five galaxies and the F1000W samples, 0.25-to-1.0 m2.1 mag for the Golden Five vs. 3.5 mag. These colors and the multi-component SED modeling discussed in the previous section indicate that the colors relative to rest 1 m are partially driven by differences in the relative luminosity of the component dominating the rest-UV (young, unobscured stellar population) and the component dominating the rest-optical (older stellar population or obscured accretion disk). As the flat UV component scales up in brightness, it leads to bluer 0.25-to-1.0 m and 0.4-to-1.0 m colors, but perhaps not because of a change in the intrinsic properties of the component dominating the optical range. It could be that a brighter UV reveals a larger fraction of the starburst emission percolating through the compact dust cloud (possibly linked to a higher burst strength or younger age, apart from dust-star relative geometry), or perhaps it shows a more massive stellar host for the obscured AGN. This interpretation also helps to explain the larger scatter in the UV region of the stacked SEDs relative to the optical region. That is, while all LRDs have distinctive blue-UV and red-optical SEDs, there is a larger diversity in the UV emission for a similar optical-to-NIR slope that might reflect variations in the relative luminosity of two different components.
@@
The overall colors of the LRDs range from 0.25-to-1.0 m 2 to 4.5 mag. Taking 0.25-to-1 m 3 mag as an intermediate value, we find that all of the Golden Five galaxies exhibit bluer colors, versus only 30% (2/7) of the galaxies in the F1000W and F770W samples. This emphasizes again that the Golden Five galaxies are intrinsically bluer than the other samples in all colors. Interestingly, we also find a trend toward stronger emission lines (larger EWs) with bluer 0.25-to-1.0 m colors. For example, this trend is seen among the bluest Golden Five sources (JADES79803 and JADES204851) and in the handful of galaxies with bluer colors in the F1000W sample (JADES210600, JADES214552 and JADES217926) and F770W sample (JADES187025 and JADES197348).
@@
The stellar masses of the Golden Five galaxies are 0.2-0.4 dex larger than those of the other subsamples, with values around for the former. Interestingly, the stellar population attenuation is lower for the Golden Five galaxies, mag, compared to F1000W sources ( mag), F770W galaxies ( mag), and non-MIRI sources ( mag).
@@ 5.4 Statistical stellar and AGN properties of LRDs
The typical stellar mass of LRDs is (median and quartiles) according to synthesizer-AGN. Considering the full redshift range of our sample of LRDs, , and the number density of galaxies in the stellar mass range detected by CEERS based on the v0.51 catalogs (Finkelstein et al., 2023), we calculate that the LRDs account for % of the full population of galaxies (subject to uncertainties due to cosmic variance). This translates to a comoving density of LRDs of Mpc-3, which is quite constant across the redshift range, with differences dex between and . The estimates for both the LRDs and the other galaxies are subject to a number of potential systematic errors, but the estimate indicates that the LRDs represent a significant, but not a dominant, population over this redshift range. Given the ranges of Universe ages probed by our sample, 200 Myr for and 400 Myr for , the % frequency within the global population could be interpreted in terms of a duty cycle around 20–40 Myr, which points to starburst behavior.
Larger masses are obtained by prospector-SF, , mainly because the mass-weighted ages are older, between 10 and 100 Myr, with median and quartiles being Myr. synthesizer-AGN fits the SEDs with significantly younger stellar populations, typically Myr. The 2 stellar populations considered by synthesizer-AGN are typically younger than 20 Myr. In fact, the average SFH of LRDs obtained by synthesizer-AGN and shown in Figure 13 indicates that LRDs are experiencing a very intense episode of star formation extending for nearly 10 Myr and with a very compact size. The burst would be in part heavily dust-enshrouded, with some younger stars having cleared the interstellar medium and being directly observable through much smaller dust optical paths. The young ages are expected for starbursts with strong emission lines (as present in some of our galaxies), with large amounts of dust (confirmed for a significant fraction of the whole sample), and with gas also feeding a SMBH. synthesizer-SF provides a similar average SFH, extended almost at a constant level up to approximately 10 Myr, and decaying afterwards. However, the first age bin considered by synthesizer-SF encompasses the 2 bursts obtained by synthesizer-AGN, the former adding more mass in ages around 10 Myr (with a large scatter, shown by the shaded region in Figure 13).
@@
The bolometric luminosities (including dust-absorbed energy) vary by a factor of 10 between the prospector-SF and prospector-AGN runs, while synthesizer-AGN lie in between, with a typical value of . Combining with the stellar masses, we conclude that LRDs present mass-to-light ratios of 1/400, typical of OB stellar associations (a B2 star would have that value, approximately), as expected based on the young ages. We remind the reader that given these very young ages and the starburst nature of LRDs, the a priori assumption of a universal IMF is quite relevant. The amount of OB stars formed, the quick and efficient formation of metals and dust, and the inferred stellar masses (or even the growth of a SMBH) are all affected by the IMF.
@@ 5.5 Spectroscopic properties of LRDs found in the literature
Apart from JADES204851, which has broad H emission at (2000 km s-1) (Matthee et al., 2023b), as mentioned in Section 5, there are other sources in our sample with relevant spectroscopic information. JADES197348 was included in the JADES NIRSpec initial data release (Bunker et al., 2023) and identified with a broad-line AGN (Maiolino et al., 2023b). Its spectrum shows a km s-1 wide component that accounts for two thirds of the H total flux, while [OIII] shows no such component. Our fits to this source are dominated by a QSO-like spectrum in the optical and NIR in the case of synthesizer-AGN and prospector-AGN. JADES154428 is found to present a broad-line component with FWHM 1800 km s-1 (Sun et al. 2024, in prep.); our fits include a non-negligible contribution from a QSO-like spectrum, dominating the SED (prospector-AGN and prospector-AGN or accounting for 50% of the emission at specific wavelengths (synthesizer-AGN). No other broad H or H line component has been reported in the remaining 18 galaxies with available spectroscopy, i.e., 17% of the spectroscopic sample are confirmed AGN hosts. For the rest of the spectroscopic sample, the presence of an AGN cannot be ruled out, since the broad-line region could be hidden due to geometrical effects. We also note that the typical depth of FRESCO NIRCam 3.9–-5.0-m grism spectroscopy is erg s-1 cm-2 for broad emission line (FWHM 1000 km s-1) from a point source and, therefore, faint broad line emission (possibly from an AGN not dominating the continuum) can remain undetected.
@@ 6 Summary and conclusions
We characterize the nature of Little Red Dots (LRDs) in the JADES field by analyzing their spectral energy distributions including the mid-infrared fluxes provided by the SMILES program for all MIRI broad-band filters. These data probe the rest-frame near-and mid-infrared where stellar emission and/or obscured AGN emission peak. After removing brown dwarfs, which contaminate our sample at the 15% level, we arrive at a sample of 31 LRDs, the surface density being 0.9 arcmin-2. This translates to a number density Mpc-3, accounting for 10–15% of the global population of galaxies with similar redshifts () and stellar mass (). Two thirds are detected in the F560W and F770W filters (all sources brighter than F444W26.5 mag), two fifths in F1000W, one seventh in F1500W, one thirteenth in F1800W, and one source in F2100W, down to 5 limits between 26.1 and 22.6 mag. The MIRI detection fraction is largely dependent on the F444W brightness, but we find an additional trend towards more detections at 10 m or beyond with bluer F277W-F444W colors.
We find that the observed MIRI colors of the LRDs, in combination with the reddest NIRCam bands, are bluer than the typical obscured QSO templates, which are dominated by the torus warm/hot dust emission at m. Indeed, the rest-frame NIR spectral range exhibits a much shallower slope that is consistent with the peak of stellar emission at around m.
@@
In general, stellar-dominated models obtain a better agreement to the near-infrared at 1-2 m as well as in the UV at m. The AGN-dominated models where emission from an obscured accretion disk dominates the optical and NIR and the torus takes over at longer wavelengths m provide a better agreement than the typical QSO templates, but still worse than the stellar models. Furthermore, this AGN-dominated model also has conceptual problems given that many of the LRDs do not present emission lines, which should be expected for the direct detection of the accretion disk (and its broad- and narrow-line regions). Consequently, we favor the interpretation that the UV-to-optical spectral range of most LRDs is dominated (% luminosity ratio) by stars.
@@
Very young stellar ages (typically around 10 Myr or younger) are supported by high equivalent width lines seen in some LRDs (with spectroscopic and imaging data), and the presence of large amounts of dust, which is a common feature of all models for the full sample, and can be expected for gas-rich, quickly enriched dense starbursts with large amounts of OB stars (affected by the IMF). However, despite the presence of large amounts of dust, LRDs are characterized by a relatively blue emission at wavelengths m, which can only be reproduced if this spectral range is dominated by (1) a QSO; (2) has holes in the interstellar medium surrounding star-forming regions that allow us to see unobscured very young star formation; or (3) have a gray attenuation law typically linked to significant scattering (with implications to the relative geometry of dust and stars and the dust clumpiness). Indeed, the large 3 mag scatter in the UV-to-NIR colors within the typical blue-red SED of the LRDs might indicate different burst strengths or different fractions of percolating light through the compact dust envelope. The presence of a QSO is hinted by some medium-band colors consistent with MgII emission. Nevertheless, for most of the sources, the best spectral energy distribution fit is obtained with just stars presenting a differentiated and/or gray dust attenuation law. A significant fraction of the total stellar emission of LRDs comes from OB stellar clumps mostly embedded in dense dusty regions with large optical paths, mag, with the integrated stellar emission obscured at the 90-95% level, which would explain the near- and mid-infrared emission jointly with some contribution from an obscured AGN.
@@
| Sample (Ngal) | redshift | code | [mag] | m-w age [Myr] | ||
@@ Appendix B Dust emission modeling
In Figure 16, we show models for a L⊙ and L⊙ compact (350 pc in size) region, attenuations mag, OB luminosity ratios 40 and 90%, and hydrogen number densities and cm-3. The templates include a fixed stellar population, which is removed for our modeling of LRDs. Models with and without stars are shown in Figure 16.
@@
The MIR slopes of the Siebenmorgen & Krügel (2007) models are very similar to the slopes of AGN torus emission presented in Nenkova et al. (2008), which are used in our prospector-AGN fits. Nenkova et al. (2008) parametrizes the emission from clumpy dust tori in terms of the number of clouds intercepting the visual, which depend on other parameters such as the torus thickness (outer to inner radii), the density and angular distribution of clumps and the viewing angle. The subset of templates included in FSPS and then Prospector adopts typical assumptions for the majority of these parameters (see e.g., Leja et al. 2018) leaving e only the scaling (overall luminosity) and the optical depth of an individual dust clump at 5500Å( from 5 to 150) as free parameters. For the prospector-AGN fits we remove any contribution from the accretion disk in the dust tori templates and we instead model the disk separately following a combination of empirical QSO template plus a power-law fν= (equivalent to fν=) with variable slope to 0.5, as discussed in Hernán-Caballero et al. 2016 or more recently in Bosman et al. 2023, attenuated by a Calzetti law. For reference, the accretion disk model with A(V)=1.5 is shown in Figure 16 in magenta. The main difference between the Nenkova et al. (2008) models and the ones in Siebenmorgen & Krügel (2007) dominated by star formation is the absence of PAH emission (especially relevant for our observations, the one at 3 m) and the silicate absorption present in some dust models. Overall, a dust torus (and an AGN template) is almost featureless.